OpenStudy (anonymous):

Solve using substitution: -y − 6x = -9 2y − 3x = -12 A. (3, -9) B. (2, -3) C. Infinitely many solutions D. No solution

OpenStudy (zzr0ck3r):

y = 9-6x then plug that intot the second equation 2(9-6x)-3x=-12 18-15x = -12 -15x=-30 x = 2 then plug that into the first -y-12=-9 y=-3 (2,-3)

OpenStudy (zzr0ck3r):

make since?